The lifetime of 225 light bulbs is measured and the sample mean and sample variance are found to be 223 hr and 100 hr2 , respectively. (a) Find a 95% confidence interval for the mean lifetime. (b) Find a 95% confidence interval for the variance of the lifetime.


1
Expert's answer
2022-02-15T00:42:02-0500

Sample mean=223hr

Sample variance=100hr2

a)

For 95%

223-2.92"\le" 223"\le" 223+5.62

220.08"\le" 223"\le" 228.62

b)

For 95%



87.3"\le" 100"\le" 112.7